The ratio of skin depth in copper at 1 kHz to that at 100 MHz is approximately :
(A) 3
(B) 30
(C) 300
(D) 3000
1
Expert's answer
2019-04-11T11:13:06-0400

Skin depth δ is given by formula


"\delta= \sqrt{\frac {\rho}{\pi f \mu}} (1)"

where ρ is the resistivity of material, f is the frequency, μ is the permeability

Using (1) we got:


"\frac {\delta_1}{\delta_2}=\sqrt{\frac {f_2}{f_1}} (2)"

In this case, we have

"\frac {\delta_1}{\delta_2}=316"

Answer:

(C) 300